OpenStudy (anonymous):

Zoe keeps track of the miles per gallon her car gets per week. She has accumulated the following data: (1, 24), (2, 24.38), (3, 24.76), (4, 25.14) What is the common difference or ratio? The common difference is 0.38. The common difference is 0.76. The common ratio is 0.99. The common ratio is 1.02.

OpenStudy (texaschic101):

25.14 - 24.76 = 0.38 24.76 - 24.38 = 0.38 24.38 - 24 = 0.38 you see the pattern.....common difference is 0.38